OpenStudy (anonymous):

find the quotient: 5y-18/-5 a. -y+3 b. -5y+15 c. -y+5

OpenStudy (anonymous):

(5y-18)/-5 or (5y)-(18/-5)

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

\[ \frac{5y-18}{-5} = \frac{5y}{-5} + \frac{18}{5} = ...?\]

OpenStudy (anonymous):

\[\frac{5y}{-5}-\frac{18}{-5}\]

OpenStudy (anonymous):

none of the answers above... Please check your question again

OpenStudy (anonymous):

i meant -15 not -18 sorry

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

\[\frac{5y-15}{-5} = \frac{5y}{-5} + \frac{15}{5} = ...?\]

OpenStudy (anonymous):

Note that 15 is divisible by 5

OpenStudy (anonymous):

So it's a

OpenStudy (anonymous):

Do you understand @tbrooks3 ?

OpenStudy (anonymous):

yes i get it thanks :)

OpenStudy (anonymous):

You're welcome!